hey i need help for my math and please also show the steps.
Thanks =)
1. 6x+7=-14y
2.4x-3/8y=1
3. y+1/3=1/4x-3
this is for the graphing linear equations but i need to only know whats x and y when they are each 0. like when x is 0 whats y and the other way around.

Answer by checkley77(12844)

For these equations set x=0 & solve for y then set y=0 & solve for x
1. 6x+7=-14y
6x+7=-14*0
6x+7=0
6x=-7
x=-7/6 (-7/6,0)ans.
6*0+7=14y
14y=7
y=7/14
y=1/2 (0,1/2) ans
-------------------------------
2.4x-3/8y=1
4x-3/8*0=1
4x=1
x=1/4 (1/4,0) ans.
4*0-3y/8=1
-3y=1*8
-3y=8
y=8/-3
y=-3/8 (0,-3/8) ans.
--------------------------------
3. y+1/3=1/4x-3
0+1/3=x/4-3
x/4=1/3+3
x/4=1/3+9/3
x/4=10/3
3x=4*10
3x=40
x=40/3 (40/3,0)
y+1/3=0/4-3
y=-3-1/3
y=-9/3-1/3
y=-10/3 (0,-10/3)